# Quantum resource degradation theory within the framework of observational entropy decomposition

### v2

@@ III.5 Results and Discussion
In Sec. IV we validate this framework by applying it to the barren-plateau problem in VQAs, and show that real-time monitoring of $\eta$ can diagnose—and potentially help mitigate—optimization stagnation.
@@
III.6 Contrast with Standard Scalar Resource Theories
@@
This distinction is particularly relevant for phenomena such as barren plateaus, where the performance of variational quantum algorithms collapses even though the underlying resource measure is conserved: in our picture, the plateau arises from a loss of resource quality, not from a loss of resource quantity.
III.7 Connection with Purity and Mixedness Measures
Conventional purity is typically quantified by the purity $\text{Tr}(\rho^{2})$ or the von Neumann entropy $S(\rho)=-\text{Tr}(\rho\log\rho)$. These are global measures characterizing a state’s deviation from the maximally mixed state. Decoherence generally increases the mixedness of a system (equivalently, decreases its purity).
@@ III.7 Connection with Purity and Mixedness Measures
III.8 Comparison with Traditional Entanglement Measures
@@
As shown in Table 3, entanglement entropy is sensitive to the overall loss of resources but cannot distinguish between the conversion of coherence into noise and a simple reduction in the total amount of resources. In contrast, $\eta$ is designed to detect resource quality degradation, providing a more refined diagnostic for optimization problems such as the barren plateau problem (BPP).
III.9 Summary of Theoretical Positioning
@@
IV Resource Degradation Theory in VQAs
The degradation mechanism we identify is not specific to VQAs; it is generic across quantum technologies. It appears in quantum sensing when phase sensitivity degrades even though the signal strength remains fixed, in quantum simulation when Trotter errors build up and act effectively as classical noise, and in quantum memories when the coherence time is limited not by leakage out of the code space but by scrambling within logical blocks.
@@ IV Resource Degradation Theory in VQAs
IV.1 Experimental Setup
IV.1.1 System Model and Hamiltonian
We study a four-qubit transverse-field Ising model with periodic boundary conditions, $Z_{5}=Z_{1}$, described by the Hamiltonian[Ising]
@@ IV.1.1 System Model and Hamiltonian
IV.1.2 Coarse-Graining Scheme
@@
IV.1.3 Cost Function
In this subsection, we define the cost function used to optimize the VQA. The cost function is the expectation value of the system Hamiltonian with respect to the parameterized quantum state,
@@ IV.1.3 Cost Function
IV.1.4 Circuit Structure
We employ a hardware-efficient ansatz[ansatz1, ansatz2], whose detailed architecture is shown in Fig. 4:
@@ IV.1.4 Circuit Structure
IV.1.5 Optimization Protocol and Parameter Update Rules
@@
IV.2 Resource evolution under degradation: theoretical and experimental analysis
This section analyzes the evolution of quantum resource metrics under degradation channels, building on the experimental framework in Sec. IV.1. We track the dynamics of the resource purity $\eta$ and its components $\mathcal{C}_{\text{rel}}$ and $\mathcal{D}_{\text{rel}}$ to demonstrate that the loss of resource quality underlies the stagnation of VQA optimization.
Figure 2: Verification of quantum resource degradation in VQAs. The figure consists of four subplots: (e) shows the evolution of the cost function, while (f) and (g) show the dynamics of the resource metrics for $\beta=0.74$. Subplot (h) shows the Pearson correlation between $\mathcal{C}_{\text{rel}}$ and $\mathcal{D}_{\text{rel}}$.
Figure 2 presents an experimental validation of quantum resource degradation through four complementary analyses.
Subplot (e) shows the progressive decline of the cost function under the three-phase degradation protocol. During Phase I (iterations 0–30), optimization proceeds efficiently in the absence of degradation. As degradation is introduced and increased in Phase II (30–60), optimization slows markedly. In Phase III (60–150), the cost function becomes effectively frozen despite continued parameter updates, indicating that the degradation of resource quality drives the onset of barren-plateau-like behavior and halts further optimization progress.
Subplot (f) illustrates the underlying transformation mechanism through the evolution of the resource metrics. Out-of-phase oscillations between $\mathcal{C}_{\text{rel}}$ and $\mathcal{D}_{\text{rel}}$ reveal a systematic conversion of quantum coherence into classical noise. During this evolution, $\mathcal{O}_{\mathcal{C}}$ remains approximately conserved, indicating a redistribution of resources rather than a net loss. The monotonic decrease of $\eta$ directly quantifies the gradual degradation of resource quality.
Subplot (g) focuses on the strong degradation regime (iterations 60–150), where the coherence-to-noise conversion is most pronounced. The persistent anticorrelation between $\mathcal{C}_{\text{rel}}$ and $\mathcal{D}_{\text{rel}}$ confirms the robustness of the conversion dynamics. Instances with $\Delta\mathcal{O}_{\mathcal{C}}\approx 0$ provide experimental support for Theorem 1, demonstrating a quantitative balance between loss of quantum coherence and generation of classical noise under conservation of the total resource.
Subplot (h) provides statistical support for this picture via a Pearson-correlation analysis.
@@ IV.2 Resource evolution under degradation: theoretical and experimental analysis
$Z=\frac{1}{2}\ln\left(\frac{1+r}{1-r}\right).$ (23)
@@
The Pearson correlation coefficient $r$ and the $p$-value in subplot (h) demonstrate a statistically significant inverse relationship. The coefficient r=0.9309r=-0.9309 quantifies the strong negative linear correlation between $\mathcal{C}_{\text{rel}}$ and $\mathcal{D}_{\text{rel}}$, while the $p$-value ($p\approx 0$) indicates extremely high statistical significance. This $p$-value represents the probability of observing such a strong correlation by random chance under the null hypothesis of no relationship. The combination of $r$ approaching 1-1 and $p<0.001$ provides compelling evidence against the null hypothesis, confirming that the observed anti-correlation is genuine and not due to random fluctuations. This statistical relationship supports the theoretical prediction of systematic conversion between coherence and noise components during resource degradation.
To further verify the statistical reliability of the negative correlation between $\mathcal{C}_{\text{rel}}$ and $\mathcal{D}_{\text{rel}}$, we performed 20 independent repetitions of the experiment under identical initial conditions (see Table 4 and Fig. 3). The Pearson correlation coefficients obtained from these runs had a mean value of 0.8999-0.8999 with a standard deviation of $0.0446$. The 95% confidence interval for the mean correlation coefficient was [0.9293,0.8590][-0.9293,-0.8590], which is well separated from zero, providing strong evidence against the null hypothesis of no correlation. All runs yielded $p$-values less than $0.001$, confirming the statistical significance of the observed anti-correlation. This consistency across multiple independent realizations demonstrates the robustness of the coherence-to-noise conversion mechanism that governs resource degradation in VQAs.
Table 4: Statistical summary of 20 independent runs with degradation channel parameter $\beta=0.74$.
@@
Figure 3: Statistical distribution of the Pearson correlation coefficients $r$ over repeated experiments. Subplot (i) shows the distribution of $r$ for $\beta=0.74$ and $N=20$. Subplot (j) shows the corresponding histogram and kernel density estimate of $r$.
@@
FIG. 4 shows the quantum circuit corresponding to the hardware-efficient ansatz described in Sec. IV.1.4. This circuit prepares the quantum states used to obtain the resource-evolution data in FIG. 2. In the final measurement stage, each qubit is mapped to a distinct classical register, $q_{0}\to c_{0}$, $q_{1}\to c_{1}$, $q_{2}\to c_{2}$, and $q_{3}\to c_{3}$. This one-to-one mapping facilitates classical post-processing and feedback within the VQA optimization loop. The circuit is designed to ensure consistent readout for evaluating the resource metrics $\eta$, $\mathcal{C}_{\text{rel}}$, and $\mathcal{D}_{\text{rel}}$ analyzed in FIG. 2.
IV.3 Resource Purity as a Diagnostic Tool
Our results indicate that the barren-plateau phenomenon (BPP) arises from a degradation in resource quality. While traditional explanations emphasize entanglement or randomness, our framework highlights the conversion of coherent resources into noise.
@@ IV.3 Resource Purity as a Diagnostic Tool
This resource-based perspective both explains the onset of BPP and informs the design of more robust variational quantum algorithms. Numerical simulations validate $\eta$ as a practical and sensitive diagnostic. Tracking resource quality in this way provides a physically transparent route toward noise-resilient algorithms and offers a cross-platform criterion for assessing quantum advantage in current and near-term noisy devices.
V Conclusion
@@
When applied to VQAs, our theory reframes optimization slowdowns as a degradation of resource quality rather than a simple depletion of resource quantity. The strong negative correlation between $\mathcal{C}_{\text{rel}}$ and $\mathcal{D}_{\text{rel}}$, demonstrated in Sec. IV.2, offers direct evidence for a coherence-to-noise conversion process. This correlation supports the presence of a systematic transformation mechanism at the core of barren plateaus.
@@
Future directions include investigating links with Loschmidt echoes, out-of-time-order correlators, and thermodynamic irreversibility, aiming to establish resource purity $\eta$ as a universal indicator of quantum resilience across different areas of physics. Further research avenues involve extending the framework to many-body and entangled resource settings; determining general conditions for degradation under free operations; constructing more refined quality metrics and efficient protocols for estimating $\eta$; realizing real-time $\eta$ tracking on noisy quantum hardware; and developing adaptive optimization schemes that actively safeguard resource quality and thus maintain quantum advantage.
@@ Acknowledgments
Appendix A: Proof of Proposition 1
@@
Appendix B: Proof of Theorem 1
@@
$\begin{aligned} \rho=&\tfrac{3}{8}|00\rangle\langle 00|+\tfrac{1}{4}|00\rangle\langle 11|+\tfrac{1}{8}|01\rangle\langle 01|\\ &+\tfrac{1}{8}|10\rangle\langle 10|+\tfrac{1}{4}|11\rangle\langle 00|+\tfrac{3}{8}|11\rangle\langle 11|.\end{aligned}$ (34)
@@
$\mathcal{C}_{\text{rel}}(\rho,\mathcal{C})=S(\Delta[\rho])-S(\rho),$ (36)
@@
$\begin{aligned} S(\rho)&=-\frac{5}{8}\log_{2}\frac{5}{8}-\frac{3}{8}\log_{2}\frac{3}{8},\\ S(\Delta[\rho])&=-\frac{3}{4}\log_{2}\frac{3}{8}-\frac{1}{4}\log_{2}\frac{1}{8}.\end{aligned}$ (38)
@@
$\begin{aligned} D[\rho_{1}\|\kappa_{1}]&=\operatorname{Tr}[\rho_{1}(\log_{2}\rho_{1}-\log_{2}\kappa_{1})]=\frac{3}{4}\log_{2}3-1,\\ D[\rho_{2}\|\kappa_{2}]&=\operatorname{Tr}[\rho_{2}(\log_{2}\rho_{2}-\log_{2}\kappa_{2})]=\frac{3}{4}\log_{2}3-1.\end{aligned}$ (41)
@@
  • Why do we choose this operation? This channel is minimal in structure yet makes the degradation mechanism transparent. With probability $\varepsilon$, the system is reset to a pure state that is block-diagonal under the chosen coarse-graining, which directly suppresses coherence between different blocks ($\Delta\mathcal{C}_{\text{rel}}<0$). At the same time, because the reset state is pure, it has maximal purity within its block and thus increases the intra-block order $\mathcal{D}_{\text{rel}}$ ($\Delta\mathcal{D}_{\text{rel}}>0$). By tuning $\varepsilon$, we can continuously control the degree of degradation and verify that $\Delta\mathcal{O}_{\mathcal{C}}\approx 0$ for suitable parameters.
@@
$\begin{aligned} \Delta[\Lambda_{\varepsilon}(\rho)]&=\hat{P}_{1}\Lambda_{\varepsilon}(\rho)\hat{P}_{1}+\hat{P}_{2}\Lambda_{\varepsilon}(\rho)\hat{P}_{2}\\ &=\frac{3+5\varepsilon}{8}\,|00\rangle\langle 00|+\frac{1-\varepsilon}{8}\,|01\rangle\langle 01|\\ &\quad+\frac{1-\varepsilon}{8}\,|10\rangle\langle 10|+\frac{3-3\varepsilon}{8}\,|11\rangle\langle 11|.\end{aligned}$ (51)
@@
$\begin{aligned} S(\Lambda_{\varepsilon}(\rho))={}&-\frac{1-\varepsilon}{4}\log_{2}\frac{1-\varepsilon}{8}-\lambda^{\varepsilon}_{3}\log_{2}\lambda^{\varepsilon}_{3}-\lambda^{\varepsilon}_{4}\log_{2}\lambda^{\varepsilon}_{4},\end{aligned}$ (52)
$\begin{aligned} S(\Delta[\Lambda_{\varepsilon}(\rho)])=&-\frac{3+5\varepsilon}{8}\log_{2}\frac{3+5\varepsilon}{8}\\ &-\frac{1-\varepsilon}{4}\log_{2}\frac{1-\varepsilon}{8}-\frac{3-3\varepsilon}{8}\log_{2}\frac{3-3\varepsilon}{8}.\end{aligned}$ (53)
@@
$\begin{aligned} \mathcal{C}_{\text{rel}}(\Lambda_{\varepsilon}(\rho),\mathcal{C})={}&-\frac{3-3\varepsilon}{8}\log_{2}\frac{3-3\varepsilon}{8}-\frac{3+5\varepsilon}{8}\log_{2}\frac{3+5\varepsilon}{8}\\ &+\lambda^{\varepsilon}_{3}\log_{2}\lambda^{\varepsilon}_{3}+\lambda^{\varepsilon}_{4}\log_{2}\lambda^{\varepsilon}_{4}.\end{aligned}$ (54)
@@
$\begin{aligned} \rho^{\varepsilon}_{1}&=\frac{\hat{P}_{1}\Lambda_{\varepsilon}(\rho)\hat{P}_{1}}{p^{\varepsilon}_{1}}=\frac{3+5\varepsilon}{4(1+\varepsilon)}\,|00\rangle\langle 00|+\frac{1-\varepsilon}{4(1+\varepsilon)}\,|01\rangle\langle 01|,\\ \rho^{\varepsilon}_{2}&=\frac{\hat{P}_{2}\Lambda_{\varepsilon}(\rho)\hat{P}_{2}}{p^{\varepsilon}_{2}}=\frac{1}{4}\,|10\rangle\langle 10|+\frac{3}{4}\,|11\rangle\langle 11|.\end{aligned}$ (55)
@@
$\begin{aligned} \mathcal{O}_{\mathcal{C}}(\Lambda_{\varepsilon}(\rho))={}&\mathcal{C}_{\text{rel}}(\Lambda_{\varepsilon}(\rho),\mathcal{C})+\mathcal{D}_{\text{rel}}(\Lambda_{\varepsilon}(\rho))\\ ={}&\lambda^{\varepsilon}_{3}\log_{2}\lambda^{\varepsilon}_{3}-\frac{3+5\varepsilon}{8}\log_{2}\frac{1+\varepsilon}{2}\\ &-\frac{3-3\varepsilon}{8}\log_{2}\frac{1-\varepsilon}{8}\\ &+\lambda^{\varepsilon}_{4}\log_{2}\lambda^{\varepsilon}_{4}+\frac{1-\varepsilon}{8}\log_{2}\frac{1-\varepsilon}{4(1+\varepsilon)}+\varepsilon.\end{aligned}$ (58)
@@
$\begin{aligned} \Delta\mathcal{C}_{\text{rel}}&=\mathcal{C}_{\text{rel}}(\Lambda_{\varepsilon}(\rho),\mathcal{C})-\mathcal{C}_{\text{rel}}(\rho,\mathcal{C})\approx-0.12986,\\ \Delta\mathcal{D}_{\text{rel}}&=\mathcal{D}_{\text{rel}}(\Lambda_{\varepsilon}(\rho))-\mathcal{D}_{\text{rel}}(\rho)\approx 0.12898,\\ \Delta\mathcal{O}_{\mathcal{C}}&=\mathcal{O}_{\mathcal{C}}(\Lambda_{\varepsilon}(\rho))-\mathcal{O}_{\mathcal{C}}(\rho)\approx-0.00089,\\ \eta(\rho)&\approx 0.58174,\quad\eta(\Lambda_{\varepsilon}(\rho))\approx 0.29450,\\ \mathcal{O}_{\mathcal{C}}(\rho)&\approx 0.45121,\quad\frac{|\Delta\mathcal{C}_{\text{rel}}|}{\mathcal{O}_{\mathcal{C}}(\rho)}\approx 0.28735.\end{aligned}$ (59)

### v3

@@ III.5 Results and Discussion
In Sec. V we validate this framework by applying it to the barren-plateau problem in VQAs, and show that real-time monitoring of $\eta$ can diagnose—and potentially help mitigate—optimization stagnation.
@@
IV Comparison with Existing Resource Theories
This section places our theory within the broader landscape of quantum resource theories. By clarifying the common foundations and distinctive features of our approach, we aim to highlight its theoretical novelty and practical utility.
IV.1 Contrast with Standard Scalar Resource Theories
@@
This distinction is particularly relevant for phenomena such as barren plateaus, where the performance of VQAs collapses even though the underlying resource measure is conserved: in our picture, the plateau arises from a loss of resource quality, not from a loss of resource quantity.
IV.2 Connection with Purity and Mixedness Measures
Conventional purity is typically quantified by the purity $\text{Tr}(\rho^{2})$ or the von Neumann entropy $S(\rho)=-\text{Tr}(\rho\log_{2}\rho)$. These are global measures characterizing a state’s deviation from the maximally mixed state. Decoherence generally increases the mixedness of a system (equivalently, decreases its purity).
@@ IV.2 Connection with Purity and Mixedness Measures
IV.3 Comparison with Traditional Entanglement Measures
@@
As shown in Table 3, entanglement entropy is sensitive to the overall loss of resources but cannot distinguish between the conversion of coherence into noise and a simple reduction in the total amount of resources. In contrast, $\eta$ is designed to detect resource quality degradation, providing a more refined diagnostic for optimization problems such as the barren plateau problem.
IV.4 Summary of Theoretical Positioning
@@
V Resource Degradation Theory in VQAs
The degradation mechanism we identify is not specific to VQAs. It is generic across quantum technologies. It appears in quantum sensing when phase sensitivity degrades even though the signal strength remains fixed, in quantum simulation when Trotter errors build up and act effectively as classical noise, and in quantum memories when the coherence time is limited not by leakage out of the code space but by scrambling within logical blocks.
@@ V Resource Degradation Theory in VQAs
V.1 Experimental Setup
V.1.1 System Model and Hamiltonian
We study a four-qubit transverse-field Ising model with periodic boundary conditions, $Z_{5}=Z_{1}$, described by the Hamiltonian[Sachdev_2011]
@@ V.1.1 System Model and Hamiltonian
V.1.2 Coarse-Graining Scheme
@@
V.1.3 Cost Function
In this subsection, we define the cost function used to optimize the VQAs. The cost function is the expectation value of the system Hamiltonian with respect to the parameterized quantum state,
@@ V.1.3 Cost Function
V.1.4 Circuit Structure
We employ a hardware-efficient ansatz[PRXQuantum.3.010313, Wang2021], whose detailed architecture is shown in Fig. 4:
@@ V.1.4 Circuit Structure
V.1.5 Optimization Protocol and Parameter Update Rules
@@
V.2 Resource evolution under degradation: theoretical and experimental analysis
This section analyzes the evolution of quantum resource metrics under degradation channels, building on the experimental framework in Sec. V.1. We track the dynamics of the resource purity $\eta$ and its components $\mathcal{C}_{\text{rel}}$ and $\mathcal{D}_{\text{rel}}$ to demonstrate that the loss of resource quality underlies the stagnation of VQA optimization.
Figure 2: Verification of quantum resource degradation in VQAs. Fig. 2(e) shows the evolution of the cost function, while Fig. 2(f) and Fig. 2(g) show the dynamics of the resource metrics for $\beta=0.74$, and Fig. 2(h) shows the Pearson correlation between $\mathcal{C}_{\text{rel}}$ and $\mathcal{D}_{\text{rel}}$.
Fig. 2 presents an experimental validation of quantum resource degradation through four complementary analyses.
Fig. 2(e) shows the progressive decline of the cost function under the three-phase degradation protocol. During Phase I (iterations 0–30), optimization proceeds efficiently in the absence of degradation. As degradation is introduced and increased in Phase II (30–60), optimization slows markedly. In Phase III (60–150), the cost function becomes effectively frozen despite continued parameter updates, indicating that the degradation of resource quality drives the onset of barren-plateau-like behavior and halts further optimization progress.
Fig. 2(f) illustrates the underlying transformation mechanism through the evolution of the resource metrics. Out-of-phase oscillations between $\mathcal{C}_{\text{rel}}$ and $\mathcal{D}_{\text{rel}}$ reveal a systematic conversion of quantum coherence into classical noise. During this evolution, $\mathcal{O}_{\mathcal{C}}$ remains approximately conserved, indicating a redistribution of resources rather than a net loss. The monotonic decrease of $\eta$ directly quantifies the gradual degradation of resource quality.
Fig. 2(g) focuses on the strong degradation regime (iterations 60–150), where the coherence-to-noise conversion is most pronounced. The persistent anticorrelation between $\mathcal{C}_{\text{rel}}$ and $\mathcal{D}_{\text{rel}}$ confirms the robustness of the conversion dynamics. Instances with $\Delta\mathcal{O}_{\mathcal{C}}\approx 0$ provide experimental support for Theorem 1, demonstrating a quantitative balance between loss of quantum coherence and generation of classical noise under conservation of the total resource.
Fig. 2(h) provides statistical support for this picture via a Pearson-correlation analysis.
@@ V.2 Resource evolution under degradation: theoretical and experimental analysis
The Pearson correlation coefficient $r$ and the $p$-value in Fig. 2(h) demonstrate a statistically significant inverse relationship. The coefficient r=0.9309r=-0.9309 quantifies the strong negative linear correlation between $\mathcal{C}_{\text{rel}}$ and $\mathcal{D}_{\text{rel}}$, while the $p$-value ($p\approx 0$) indicates extremely high statistical significance. This $p$-value represents the probability of observing such a strong correlation by random chance under the null hypothesis of no relationship. The combination of $r$ approaching 1-1 and $p<0.001$ provides compelling evidence against the null hypothesis, confirming that the observed anti-correlation is genuine and not due to random fluctuations. This statistical relationship supports the theoretical prediction of systematic conversion between coherence and noise components during resource degradation.
@@
$Z=\frac{1}{2}\log_{2}\left(\frac{1+r}{1-r}\right).$ (23)
@@
To further verify the statistical reliability of the negative correlation between $\mathcal{C}_{\text{rel}}$ and $\mathcal{D}_{\text{rel}}$, we have performed 20 independent repetitions of the experiment under identical initial conditions (see Table 4 and Fig. 3). The Pearson correlation coefficients obtained from these runs have a mean value of 0.8999-0.8999 with a standard deviation of $0.0446$. The 95% confidence interval for the mean correlation coefficient is [0.9293,0.8590][-0.9293,-0.8590], which is well separated from zero, providing strong evidence against the null hypothesis of no correlation. All runs yielded $p$-values less than $0.001$, confirming the statistical significance of the observed anti-correlation. This consistency across multiple independent realizations demonstrates the robustness of the coherence-to-noise conversion mechanism that governs resource degradation in VQAs.
Table 4: Statistical summary of 20 independent runs with $\beta=0.74$.
@@
Figure 3: Statistical distribution of the Pearson correlation coefficients $r$ over repeated experiments. Fig. 3(i) shows a combined box plot and violin plot of the correlation coefficients. The light blue box represents the interquartile range (IQR: -0.9361 to -0.8718), containing the middle 50% of the data. The dark red line inside the box indicates the median value (-0.9108). The light green violin shape illustrates the probability density distribution of the coefficients. The red square marks the mean value (-0.8999). The gray shaded region with dashed boundaries indicates the 95% confidence interval [0.9293,0.8590][-0.9293,-0.8590]. Fig. 3(j) shows a histogram with kernel density estimation (KDE).The blue bars show the frequency distribution of correlation coefficients. The dark blue curve represents the KDE, providing a smooth estimate of the probability density. The solid red vertical line indicates the mean value (-0.8999), while the dashed gray lines and gray shaded area mark the 95% confidence interval.
The consistent negative correlation observed across all experiments (mean r=0.8999r=-0.8999, $95\%$ CI [0.9293,0.8590][-0.9293,-0.8590]) confirms the robustness of the coherence-to-noise transformation mechanism. The narrow confidence interval and small standard deviation ($0.0446$) indicate high reproducibility of the results. The interquartile range of 0.0643 demonstrates moderate variability in the correlation coefficients across experiments.
@@
Fig. 4 shows the quantum circuit corresponding to the hardware-efficient ansatz described in Sec. V.1.4. This circuit prepares the quantum states used to obtain the resource-evolution data in Fig. 2. In the final measurement stage, each qubit is mapped to a distinct classical register, $q_{0}\to c_{0}$, $q_{1}\to c_{1}$, $q_{2}\to c_{2}$, and $q_{3}\to c_{3}$. This one-to-one mapping facilitates classical post-processing and feedback within the VQA optimization loop. The circuit is designed to ensure consistent readout for evaluating the resource metrics $\eta$, $\mathcal{C}_{\text{rel}}$, and $\mathcal{D}_{\text{rel}}$ analyzed in Fig. 2.
V.3 Resource Purity as a Diagnostic Tool
Our results indicate that the BPP arises from a degradation in resource quality. While traditional explanations emphasize entanglement or randomness, our framework highlights the conversion of coherent resources into noise.
@@ V.3 Resource Purity as a Diagnostic Tool
This resource-based perspective both explains the onset of BPP and informs the design of more robust VQAs. Numerical simulations validate $\eta$ as a practical and sensitive diagnostic. Tracking resource quality in this way provides a physically transparent route toward noise-resilient algorithms and offers a cross-platform criterion for assessing quantum advantage in current and near-term noisy devices.
VI Conclusion
@@
When applied to VQAs, our theory reframes optimization slowdowns as a degradation of resource quality rather than a simple depletion of resource quantity. The strong negative correlation between $\mathcal{C}_{\text{rel}}$ and $\mathcal{D}_{\text{rel}}$, demonstrated in Sec. V.2, offers direct evidence for a coherence-to-noise conversion process. This correlation supports the presence of a systematic transformation mechanism at the core of barren plateaus.
@@
Future work can be approached at several levels.
On the application side, one may investigate how the $\eta$ indicator can enable early warnings and adaptive optimization in variational quantum algorithms.
Theoretically, it would be interesting to explore connections between this framework and dynamical signatures such as the Loschmidt echo, and to examine possible extensions to many-body entangled systems.
On the practical side, developing lightweight protocols for estimating $\eta$ on near-term noisy quantum devices would be a valuable next step.
@@ Acknowledgments
VII Appendix A: Proof of Proposition 1
@@
VIII Appendix B: Proof of Theorem 1
@@
$\begin{aligned} \rho&=\tfrac{3}{8}|00\rangle\langle 00|+\tfrac{1}{4}|00\rangle\langle 11|+\tfrac{1}{8}|01\rangle\langle 01|\\ &+\tfrac{1}{8}|10\rangle\langle 10|+\tfrac{1}{4}|11\rangle\langle 00|+\tfrac{3}{8}|11\rangle\langle 11|.\end{aligned}$ (34)
@@
$\mathcal{C}_{\text{rel}}(\rho,\mathcal{C})=S(\Delta(\rho))-S(\rho),$ (36)
@@
$\begin{aligned} S(\rho)&=-\frac{5}{8}\log_{2}\frac{5}{8}-\frac{3}{8}\log_{2}\frac{3}{8},\\ S(\Delta(\rho))&=-\frac{3}{4}\log_{2}\frac{3}{8}-\frac{1}{4}\log_{2}\frac{1}{8}.\end{aligned}$ (38)
@@
$\begin{aligned} D[\rho_{1}\|\kappa_{1}]&=\operatorname{Tr}[\rho_{1}(\log_{2}\rho_{1}-\log_{2}\kappa_{1})]\\ &=\frac{3}{4}\log_{2}3-1,\\ D[\rho_{2}\|\kappa_{2}]&=\operatorname{Tr}[\rho_{2}(\log_{2}\rho_{2}-\log_{2}\kappa_{2})]\\ &=\frac{3}{4}\log_{2}3-1.\end{aligned}$ (41)
@@
  • Why do we choose this operation? This channel is minimal in structure, yet making the degradation mechanism transparent. With probability $\varepsilon$, the system is reset to a pure state that is block-diagonal under the chosen coarse-graining, which directly suppresses coherence between different blocks ($\Delta\mathcal{C}_{\text{rel}}<0$). At the same time, because the reset state is pure, it has maximal purity within its block and thus increases the intra-block order $\mathcal{D}_{\text{rel}}$ ($\Delta\mathcal{D}_{\text{rel}}>0$). By tuning $\varepsilon$, we can continuously control the degree of degradation and verify that $\Delta\mathcal{O}_{\mathcal{C}}\approx 0$ for suitable parameters.
@@
$\begin{aligned} \Delta(\Lambda_{\varepsilon}(\rho))&=\hat{P}_{1}\Lambda_{\varepsilon}(\rho)\hat{P}_{1}+\hat{P}_{2}\Lambda_{\varepsilon}(\rho)\hat{P}_{2}\\ &=\frac{3+5\varepsilon}{8}\,|00\rangle\langle 00|+\frac{1-\varepsilon}{8}\,|01\rangle\langle 01|\\ &\quad+\frac{1-\varepsilon}{8}\,|10\rangle\langle 10|+\frac{3-3\varepsilon}{8}\,|11\rangle\langle 11|.\end{aligned}$ (51)
@@
$\begin{aligned} S(\Lambda_{\varepsilon}(\rho))=&-\frac{1-\varepsilon}{4}\log_{2}\frac{1-\varepsilon}{8}\\ &-\lambda^{\varepsilon}_{3}\log_{2}\lambda^{\varepsilon}_{3}-\lambda^{\varepsilon}_{4}\log_{2}\lambda^{\varepsilon}_{4},\end{aligned}$ (52)
$\begin{aligned} S(\Delta(\Lambda_{\varepsilon}(\rho)))=&-\frac{3+5\varepsilon}{8}\log_{2}\frac{3+5\varepsilon}{8}\\ &-\frac{1-\varepsilon}{4}\log_{2}\frac{1-\varepsilon}{8}\\ &-\frac{3-3\varepsilon}{8}\log_{2}\frac{3-3\varepsilon}{8}.\end{aligned}$ (53)
@@
$\begin{aligned} \mathcal{C}_{\text{rel}}(\Lambda_{\varepsilon}(\rho),\mathcal{C})=&-\frac{3-3\varepsilon}{8}\log_{2}\frac{3-3\varepsilon}{8}\\ &-\frac{3+5\varepsilon}{8}\log_{2}\frac{3+5\varepsilon}{8}\\ &+\lambda^{\varepsilon}_{3}\log_{2}\lambda^{\varepsilon}_{3}+\lambda^{\varepsilon}_{4}\log_{2}\lambda^{\varepsilon}_{4}.\end{aligned}$ (54)
@@
$\begin{aligned} \rho^{\varepsilon}_{1}=\frac{\hat{P}_{1}\Lambda_{\varepsilon}(\rho)\hat{P}_{1}}{p^{\varepsilon}_{1}}=&\frac{3+5\varepsilon}{4(1+\varepsilon)}\,|00\rangle\langle 00|\\ &+\frac{1-\varepsilon}{4(1+\varepsilon)}\,|01\rangle\langle 01|,\end{aligned}$ (55)
$\rho^{\varepsilon}_{2}=\frac{\hat{P}_{2}\Lambda_{\varepsilon}(\rho)\hat{P}_{2}}{p^{\varepsilon}_{2}}=\frac{1}{4}\,|10\rangle\langle 10|+\frac{3}{4}\,|11\rangle\langle 11|.$ (56)
@@
$\begin{aligned} \mathcal{O}_{\mathcal{C}}(\Lambda_{\varepsilon}(\rho))={}&\mathcal{C}_{\text{rel}}(\Lambda_{\varepsilon}(\rho),\mathcal{C})+\mathcal{D}_{\text{rel}}(\Lambda_{\varepsilon}(\rho))\\ ={}&\lambda^{\varepsilon}_{3}\log_{2}\lambda^{\varepsilon}_{3}-\frac{3+5\varepsilon}{8}\log_{2}\frac{1+\varepsilon}{2}\\ &-\frac{3-3\varepsilon}{8}\log_{2}\frac{1-\varepsilon}{8}+\lambda^{\varepsilon}_{4}\log_{2}\lambda^{\varepsilon}_{4}\\ &+\frac{1-\varepsilon}{8}\log_{2}\frac{1-\varepsilon}{4(1+\varepsilon)}+\varepsilon.\end{aligned}$ (59)
@@
$\begin{aligned} \Delta\mathcal{C}_{\text{rel}}&=\mathcal{C}_{\text{rel}}(\Lambda_{\varepsilon}(\rho),\mathcal{C})-\mathcal{C}_{\text{rel}}(\rho,\mathcal{C})\approx-0.12986,\\ \Delta\mathcal{D}_{\text{rel}}&=\mathcal{D}_{\text{rel}}(\Lambda_{\varepsilon}(\rho))-\mathcal{D}_{\text{rel}}(\rho)\approx 0.12898,\\ \Delta\mathcal{O}_{\mathcal{C}}&=\mathcal{O}_{\mathcal{C}}(\Lambda_{\varepsilon}(\rho))-\mathcal{O}_{\mathcal{C}}(\rho)\approx-0.00089,\end{aligned}$ (60)
and
$\begin{aligned} \eta(\rho)&\approx 0.58174,\quad\eta(\Lambda_{\varepsilon}(\rho))\approx 0.29450,\\ \mathcal{O}_{\mathcal{C}}(\rho)&\approx 0.45121,\quad\frac{|\Delta\mathcal{C}_{\text{rel}}|}{\mathcal{O}_{\mathcal{C}}(\rho)}\approx 0.28735.\end{aligned}$ (61)